OpenStudy (anonymous):

Solve x2 – 4x = 12 by completing the square. Show your work for full credit.

OpenStudy (cwrw238):

first you divide coefficient of -4x by 2 giving -2 and now you write (x - 2)^2 - (-2)^2 = 12 (x - 2)^2 - 4 = 12 (x - 2)^2 = 16 take square roots of both sides x - 2 = +/- 4 x = 4 +2 or -4 + 2 so the solution set is {6 , -2}

OpenStudy (anonymous):

thankyou!

OpenStudy (cwrw238):

yw